Circular No. 9077

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET P/2009 S4 (CATALINA-LINEAR) J. Scotti, Lunar and Planetary Laboratory, reports his recovery of

comet P/2004 EW_{38} (cf. IAUC 8322), noting only stellar appearance, on CCD images taken with the 1.8-m Spacewatch telescope at Kitt Peak.

2009 UT  α_{2000}  δ_{2000}  Mag.
Sept.21.37293  2^{h}47^{m}29.^{s}46  +13^{o}09′41′′.3  22.0
23.33491  2 46 55.91  +13 08 18.5
28.44718  2 45 04.37  +13 03 17.8  21.9

The indicated correction to the orbit prediction on MPC 59600 is ΔT = +0.02 day. Additional astrometry, the following linked orbital elements, residuals, and an ephemeris appear on MPEC 2009-S126.

Epoch = 2010 Sept. 1.0 TT

T = 2010 Sept. 3.7086 TT  ω = 90.^{o}1359
e = 0.500013  Ω = 49.8835 }2000.0
q = 1.794822 AU  i = 6.5249

a = 3.589741 AU n^{o} = 0.1449137 P = 6.80 years

COMETS C/2009 E2–E5, C/2009 F8, AND C/2009 F9 (SOHO) Further to IAUCs 9065 and 9070, additional presumed comets have

been found on SOHO website images — all Kreutz sungrazers except for C/2009 E3 (Meyer group). C/2009 E2 (also found by B. Zhou) was small and slightly diffuse (mag ∼ 7). C/2009 E3 was tiny and stellar in appearance (mag ∼ 7.5–8), and it faded rapidly. C/2009 E4 was small and slightly diffuse (mag ∼ 6.5–7) in C3 images; in C2 images it was faint, diffuse, and slightly elongated. C/2009 E5 was very diffuse (mag ∼ 8). C/2009 F8 was slightly diffuse (mag ∼ 7–7.5) in C3 images, and small and elongated in C2 images. C/2009 F9 was very small and condensed (mag ∼ 7.5) in C3 images, and slightly elongated (but stellar in appearance) in C2 images.

Comet 2009 UT α_{2000} δ_{2000} Inst. F MPEC
C/2009 E2 Mar. 1.221 23^{h}07.^{m}6 − 8^{o}16′ C3  MK 2009-N01
C/2009 E3 7.463 23 17.1 − 3 27 C2 EB 2009-N01
C/2009 E4  10.138 23 44.8 − 5 09 C3/2 WX 2009-N01
C/2009 E5 10.229 23 30.5 − 4 49 C2 BZ 2009-N29
C/2009 F8  16.862 0 04.1 − 2 09 C3/2 MU 2009-N29
C/2009 F9  28.471 0 44.4 + 2 17 C3/2 HS 2009-N29
2009 September 28  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2009 CBAT  Daniel W. E. Green